OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 2 and contains the point (4, 2).

OpenStudy (anonymous):

the Eqn. of a line = > y -y1 = m(x-x1) => y-2 = 2 (x-4) => y-2 = 2x - 8 =>-2x+y = -8+2 => -2x +y = -6 => 2x-y = 6